# Efficient Optimization of Hierarchical Identifiers for Generative Recommendation

### v1

@@ 1 Introduction
Traditional retrieval models address this challenge through dual-encoder architectures, which encode users and items into a shared representation space. Retrieval is then performed through (approximate) nearest neighbor algorithms over these embeddings. This approach relies on the simplifying assumption that user–item relevance can be captured by an inner product in the embedding space. Previous research argues that this assumption often fails, leading to degraded recommendation quality [comirec].
Generative retrieval models have recently emerged as a promising alternative. Framing retrieval as a sequence generation task offers greater flexibility in how users and items are represented and matched. However, these models typically face two important limitations: (i) inference efficiency at a large scale, and (ii) retrieval quality degradation due to insufficient exploitation of the compositional semantics of items [10.1145/2959100.2959190].
To address these challenges of inference efficiency and retrieval quality, Si et al. introduced SEATER [seater]. SEATER improves inference efficiency by mapping items to balanced, tree-structured identifiers, while preserving retrieval quality through contrastive training objectives that enforce semantic consistency across the tree. Their results show that their identifier design can both improve inference latency and retrieval quality.
While SEATER improves inference efficiency, its tree-construction algorithm introduces a substantial training bottleneck, especially as the item catalog grows. Our experimental results confirm this, with tree construction time taking up to 19.4% of total training time on the largest dataset (Table 5). In large-scale, frequently updated recommendation settings, this step becomes a key limitation, as tree construction must often be repeated when new items or user interactions appear. Generative retrieval models are particularly sensitive to this issue, since the retrieval index is encoded directly in model parameters—making fast retraining essential.
@@
We verify that SEATER’s tree-construction algorithm indeed improves retrieval quality over strong baselines, including the new Yambda dataset (RQ1). We then demonstrate that a greedy clustering method massively reduces tree construction time, to less than 2% of the original on the largest dataset, Yambda, while causing only a minor drop in retrieval quality (RQ2). Finally, we show that a hybrid method, combining greedy clustering at higher tree levels with full constrained clustering at lower levels, achieves retrieval quality on par with or even exceeding the original, while cutting construction time to just 5–8% across datasets (RQ3). We publish all of our code for full reproducibility.111https://github.com/joshrosie/re-seater
@@ 2.0.1 Retrieval for recommender systems
Traditional retrieval models for recommendations use dual-encoder architectures to represent users and items as embeddings. Efficient (approximate) nearest-neighbor algorithms then perform matching between user and item representations [10.1145/2959100.2959190]. Models such as GRU4Rec [gru4rec], SASRec [sasrec], and BERT4Rec [bert4rec] predict the next item from user interaction histories, typically treating item IDs as discrete tokens and using recurrent or transformer-based architectures to capture temporal patterns. Alternatively, tree-based methods like TDM [tdm18] and JTM [jtm20] improve retrieval efficiency by hierarchically organizing items for faster candidate generation. SEATER builds on these tree-based approaches by learning to optimize item identifiers from an indexing perspective [seater].
@@ 2.0.2 Generative retrieval
Generative retrieval frames retrieval as sequence generation over discrete identifiers, and was first explored in document search with models such as DSI and NCI [10.1145/3476415.3476428, DBLP:conf/nips/Tay00NBM000GSCM22, 10.1145/3626772.3657697]. In recommendation, TIGER [rajput2023recommender] adapts this paradigm by learning semantic identifiers for items via vector quantization and training an encoder–decoder model to generate the corresponding code sequences. While effective, TIGER’s identifiers are not explicitly constrained to follow a structured hierarchy and therefore do not optimize structural properties during training. SEATER extends this paradigm by learning balanced, tree-structured identifiers and introducing contrastive objectives that align the semantic and structural properties of the learned tree. Decoding is performed through constrained beam search restricted to valid tree paths, making the retrieval process explicitly index-aware.
@@ 2.0.3 Document identifier construction
A key component of generative retrieval is the design of discrete item or document identifiers. Early work in this area used vector quantization and discrete code learning to represent semantic content compactly [pq, semantic-hashing, vqvae]. In recommendation, models such as TIGER [rajput2023recommender] adopt this idea by mapping item embeddings to sequences of quantized codes, allowing an encoder–decoder model to autoregressively generate identifiers. While this captures semantic similarity, these identifiers are not explicitly optimized for structural consistency, resulting in uneven semantic granularity across items.
SEATER takes a different approach by explicitly constructing balanced, hierarchical identifiers. Items are organized into a -ary tree through constrained clustering, ensuring equal-length identifiers where tokens represent increasingly fine-grained semantic partitions. By coupling this construction with contrastive learning to align item semantics with the tree structure [seater], SEATER produces identifiers that combine semantic coherence with structural regularity, yielding faster inference and improved retrieval quality.
@@ 3 Methodology
We first summarize the SEATER model [seater], which we reproduce as the basis for our study. We then present our proposed modifications to its tree construction algorithm to improve training efficiency while maintaining inference efficiency and retrieval quality.
@@ 3.1 SEATER
Given a user’s interaction history, SEATER generates an item identifier: a sequence of tokens that uniquely represents the item to be retrieved. The model uses a lightweight transformer encoder–decoder architecture to perform this task. Each item in the catalog is assigned a unique identifier composed of multiple tokens, corresponding to nodes in a pre-constructed balanced -ary tree. During retrieval, SEATER generates these identifiers using constrained beam search, ensuring that decoding follows valid token paths within the tree.
@@ 3.1.1 Item identifiers
The balanced tree structure reduces search complexity from in naive decoding to , where is the beam width and the number of items. This leads to substantial inference-time efficiency gains over prior generative retrieval models [seater], as summarized in Table 1.
@@ 3.2 Tree construction algorithms
which scales poorly as increases [Levy-Kramer_k-means-constrained_2018]. Moreover, because the capacity-constrained assignment step couples all item–cluster decisions, it is difficult to parallelize, limiting scalability on modern GPU hardware. As shown in 1(a), this constrained -means solution towards a balanced tree construction restricts SEATER’s practicality for large datasets or frequent retraining.
@@ 3.2.2 Hybrid clustering
In summary, our greedy and hybrid clustering algorithms address SEATER’s scalability bottleneck by replacing the global capacity-constrained optimization with a parallelizable local heuristic and combining it with precise clustering near the tree leaves. These methods reduce tree construction time by up to two orders of magnitude—down to less than 2% of the original on the largest dataset, while preserving inference efficiency.
@@ 4.1 Datasets
Following the original work [seater], we evaluate SEATER on Yelp, Books, and News, and extend the evaluation to a larger-scale dataset, Yambda. Table 2 summarizes the main statistics.
@@ 4.1.1 Yelp
Yelp contains user–business interactions and reviews.222https://business.yelp.com/data/resources/open-dataset/ We use the standard train, validation, and test splits provided in the LightGCN repository [10.1145/3397271.3401063]. This setup was verified with the SEATER authors to align with their original preprocessing.
@@ 4.1.2 Books
The Books subset of the Amazon review dataset [ni-etal-2019-justifying].333https://nijianmo.github.io/amazon/ We follow the ComiRec preprocessing pipeline [comirec], identical to the one used in SEATER’s original implementation.
@@ 4.1.3 News
The small partition of the MIcrosoft News Dataset (MIND) [wu-etal-2020-mind].444https://msnews.github.io/ Since the original preprocessing code was unavailable, we reconstructed the setup based on author correspondence. We used the training data of MIND-small and split users by an 8:1:1 ratio into train, validation, and test sets. For each user, positive impressions were concatenated with their interaction history, and sequences were expanded into multiple training instances. During evaluation, we used the first 80% of each sequence as input context and the remaining 20% as targets, consistent with the other datasets.
@@ 4.1.4 Yambda
Yambda [ploshkin2025yambda5blargescalemultimodal] is a large-scale open dataset of user–music interactions from Yandex Music.555https://huggingface.co/datasets/yandex/yambda With a substantially larger item pool, it serves as a scalability stress test, clearly exposing tree construction as a bottleneck during training and motivating our proposed algorithms. Following the dataset creators’ preprocessing, we keep only organic interactions and filter out tracks played for less than 50% of their duration. Data is temporally split into 299 days for training, a 30-minute gap, one day for validation, and one day for testing.
@@ 4.2.1 Baselines
We compare SEATER against the baselines reported in the original paper, using the authors’ published results. To verify reproducibility and ensure consistent evaluation, we re-implement three representative baselines: GRU4Rec, SASRec, and BERT4Rec, following the configurations described in the original work [seater]. GRU4Rec is implemented via the ComiRec repository.777https://github.com/THUDM/ComiRec/tree/master We implement BERT4Rec using an open-source implementation888https://github.com/jaywonchung/BERT4Rec-VAE-Pytorch, and SASRec based on the SEATER authors’ code. GRU4Rec and SASRec are trained autoregressively by predicting the next item from user history, while BERT4Rec uses a bidirectional Cloze-style objective. Learning rates are selected from {0.001, 0.0001}, with 0.001 performing best in most cases and 0.0001 used for BERT4Rec on the News dataset.
@@ 4.2.2 Tree construction
We build SEATER’s identifier tree using SASRec embeddings trained on the same dataset. These embeddings are only used during tree construction, and then discarded. We use the tree construction implementation provided by the original authors [seater], which uses constrained -means [Levy-Kramer_k-means-constrained_2018]. Our greedy variant replaces it with a Faiss-based approximate nearest-neighbor search [douze2024faiss], and the hybrid method switches from greedy to constrained clustering once the number of items per level drops below 2,000, an empirically chosen threshold balancing efficiency and clustering precision.
@@ 4.2.3 Evaluation protocol
Following [seater], we evaluate retrieval accuracy using Recall@k, Hit Rate@k (HR@k), and Normalized Discounted Cumulative Gain@k (NDCG@k) for , averaged across test users. To assess training efficiency, we measure total training time (excluding tree construction) and tree-construction time. This allows us to isolate the cost of tree building and highlight its impact as a potential training time bottleneck. It is worth noting that, next to our reported tree-construction and training times, the full SEATER pipeline also includes training SASRec to obtain item embeddings. Following the original work [seater], we assume that practitioners already have access to an existing embedding model and therefore exclude SASRec training time from our comparisons.
@@ 5.2 RQ2: The impact of greedy tree construction
Using greedy tree construction in SEATER substantially reduces tree-building time across all datasets, as shown in Table 5 and 1(b). The improvement is particularly pronounced for datasets with larger item catalogs, such as Books and Yambda, where tree-construction time decreases from 54 minutes to 93 seconds and from 82 minutes to 83 seconds, respectively—representing a 35 and 59 speedup. Interestingly, even when excluding tree construction, overall training time also decreases, likely due to faster model convergence. We leave it to future work to investigate the relationship between item identifier structure and convergence dynamics. Despite the use of an approximate (greedy) clustering algorithm, the performance drop across all metrics remains minimal, with the largest observed difference being only in Recall@20 on the Books dataset. To answer RQ2, the proposed greedy clustering method offers a substantial improvement in training efficiency with a minimal drop in retrieval quality.
@@ 5.3 RQ3: Balancing greedy and complete clustering
| News | NDCG@20 | 0.0916 | 0.0959 | 0.0950 | 0.0957 |
| NDCG@50 | 0.1209 | 0.1251 | 0.1246 | 0.1249 | |
| Recall@20 | 0.1516 | 0.1581 | 0.1548 | 0.1552 | |
| Recall@50 | 0.2524 | 0.2581 | 0.2576 | 0.2555 | |
| Hit@20 | 0.4106 | 0.4236 | 0.4151 | 0.4155 | |
| Hit@50 | 0.5792 | 0.5810 | 0.5812 | 0.5755 | |
| Tree time (s) | – | 190.02 | 67.08 | 31.61 | |
| Training time (h:m:s) | 0:11:13 | 1:37:48 | 1:37:18 | 1:35:56 | |
| Yelp | NDCG@20 | 0.0468 | 0.0527 | 0.0515 | 0.0500 |
| NDCG@50 | 0.0694 | 0.0756 | 0.0742 | 0.0720 | |
| Recall@20 | 0.0597 | 0.0654 | 0.0639 | 0.0616 | |
| Recall@50 | 0.1195 | 0.1268 | 0.1249 | 0.1208 | |
| Hit@20 | 0.3762 | 0.3974 | 0.3966 | 0.3844 | |
| Hit@50 | 0.5859 | 0.5940 | 0.5900 | 0.5791 | |
| Tree time (s) | – | 189.35 | 61.27 | 31.15 | |
| Training time (h:m:s) | 0:18:42 | 1:10:33 | 1:10:14 | 0:59:42 | |
| Books | NDCG@20 | 0.0408 | 0.0593 | 0.0574 | 0.0566 |
| NDCG@50 | 0.0536 | 0.0714 | 0.0691 | 0.0682 | |
| Recall@20 | 0.0807 | 0.0967 | 0.0936 | 0.0922 | |
| Recall@50 | 0.1311 | 0.1442 | 0.1398 | 0.1374 | |
| Hit@20 | 0.1689 | 0.2000 | 0.1960 | 0.1931 | |
| Hit@50 | 0.2578 | 0.2815 | 0.2756 | 0.2720 | |
| Tree time (s) | – | 3232.80 | 184.05 | 93.32 | |
| Training time (h:m:s) | 0:26:49 | 3:54:27 | 3:04:38 | 2:28:32 | |
| Yambda | NDCG@20 | 0.0960 | 0.1245 | 0.1495 | 0.1399 |
| NDCG@50 | 0.1138 | 0.1370 | 0.1601 | 0.1493 | |
| Recall@20 | 0.1006 | 0.1175 | 0.1461 | 0.1353 | |
| Recall@50 | 0.1636 | 0.1705 | 0.1966 | 0.1817 | |
| Hit@20 | 0.4298 | 0.4566 | 0.4975 | 0.4841 | |
| Hit@50 | 0.5406 | 0.5485 | 0.5795 | 0.5638 | |
| Tree time (s) | – | 4907.97 | 375.72 | 83.21 | |
| Training time (h:m:s) | 0:47:16 | 5:39:04 | 4:23:58 | 2:37:05 | |

### v2

@@ 1 Introduction
Traditional retrieval models address this challenge through dual-encoder architectures, which encode users and items into a shared representation space. Retrieval is then performed through (approximate) nearest neighbor algorithms over these embeddings. This approach relies on the simplifying assumption that user-item relevance can be captured by an inner product in the embedding space. Previous research argues that this assumption often fails, leading to degraded recommendation quality [1].
Generative retrieval models have recently emerged as a promising alternative. Framing retrieval as a sequence generation task offers greater flexibility in how users and items are represented and matched. However, these models typically face two important limitations: (i) inference efficiency at a large scale, and (ii) retrieval quality degradation due to insufficient exploitation of the compositional semantics of items [2].
To address these challenges of inference efficiency and retrieval quality, Si et al. introduced SEATER [14]. SEATER improves inference efficiency by mapping items to balanced, tree-structured identifiers, while preserving retrieval quality through contrastive training objectives that enforce semantic consistency across the tree. Their results show that their identifier design can both improve inference latency and retrieval quality.
While SEATER improves inference efficiency, its tree-construction algorithm introduces a substantial training bottleneck, especially as the item catalog grows. Our experimental results confirm this, with tree construction time taking up to 19.4% of total training time on the largest dataset (Table 5). In large-scale, frequently updated recommendation settings, this step becomes a key limitation, as tree construction must often be repeated when new items or user interactions appear. Generative retrieval models are particularly sensitive to this issue, since the retrieval index is encoded directly in model parameters, making fast retraining essential.
@@
We verify that SEATER’s tree-construction algorithm indeed improves retrieval quality over strong baselines, including the new Yambda dataset (RQ1). We then demonstrate that a greedy clustering method massively reduces tree construction time, to less than 2% of the original on the largest dataset, Yambda, while causing only a minor drop in retrieval quality (RQ2). Finally, we show that a hybrid method, combining greedy clustering at higher tree levels with full constrained clustering at lower levels, achieves retrieval quality on par with or even exceeding the original, while cutting construction time to just 5-8% across datasets (RQ3). We publish all of our code for full reproducibility.111https://github.com/joshrosie/re-seater
@@ 2.0.1 Retrieval for recommender systems
Traditional retrieval models for recommendations use dual-encoder architectures to represent users and items as embeddings. Efficient (approximate) nearest-neighbor algorithms then perform matching between user and item representations [2]. Models such as GRU4Rec [5], SASRec [7], and BERT4Rec [15] predict the next item from user interaction histories, typically treating item IDs as discrete tokens and using recurrent or transformer-based architectures to capture temporal patterns. Alternatively, tree-based methods like TDM [22] and JTM [21] improve retrieval efficiency by hierarchically organizing items for faster candidate generation. SEATER builds on these tree-based approaches by learning to optimize item identifiers from an indexing perspective [14].
@@ 2.0.2 Generative retrieval
Generative retrieval frames retrieval as sequence generation over discrete identifiers, and was first explored in document search with models such as DSI and NCI [9, 16, 20]. In recommendation, TIGER [12] adapts this paradigm by learning semantic identifiers for items via vector quantization and training an encoder-decoder model to generate the corresponding code sequences. While effective, TIGER’s identifiers are not explicitly constrained to follow a structured hierarchy and therefore do not optimize structural properties during training. SEATER extends this paradigm by learning balanced, tree-structured identifiers and introducing contrastive objectives that align the semantic and structural properties of the learned tree. Decoding is performed through constrained beam search restricted to valid tree paths, making the retrieval process explicitly index-aware.
@@ 2.0.3 Document identifier construction
A key component of generative retrieval is the design of discrete item or document identifiers. Early work in this area used vector quantization and discrete code learning to represent semantic content compactly [6, 13, 17]. In recommendation, models such as TIGER [12] adopt this idea by mapping item embeddings to sequences of quantized codes, allowing an encoder-decoder model to autoregressively generate identifiers. While this captures semantic similarity, these identifiers are not explicitly optimized for structural consistency, resulting in uneven semantic granularity across items.
SEATER takes a different approach by explicitly constructing balanced, hierarchical identifiers. Items are organized into a -ary tree through constrained clustering, ensuring equal-length identifiers where tokens represent increasingly fine-grained semantic partitions. By coupling this construction with contrastive learning to align item semantics with the tree structure [14], SEATER produces identifiers that combine semantic coherence with structural regularity, yielding faster inference and improved retrieval quality.
@@ 3 Methodology
We first summarize the SEATER model [14], which we reproduce as the basis for our study. We then present our proposed modifications to its tree construction algorithm to improve training efficiency while maintaining inference efficiency and retrieval quality.
@@ 3.1 SEATER
Given a user’s interaction history, SEATER generates an item identifier: a sequence of tokens that uniquely represents the item to be retrieved. The model uses a lightweight transformer encoder-decoder architecture to perform this task. Each item in the catalog is assigned a unique identifier composed of multiple tokens, corresponding to nodes in a pre-constructed balanced -ary tree. During retrieval, SEATER generates these identifiers using constrained beam search, ensuring that decoding follows valid token paths within the tree.
@@ 3.1.1 Item identifiers
The balanced tree structure reduces search complexity from in naive decoding to , where is the beam width and the number of items. This leads to substantial inference-time efficiency gains over prior generative retrieval models [14], as summarized in Table 1.
@@ 3.2 Tree construction algorithms
which scales poorly as increases [8]. Moreover, because the capacity-constrained assignment step couples all item-cluster decisions, it is difficult to parallelize, limiting scalability on modern GPU hardware. As shown in 1(a), this constrained -means solution towards a balanced tree construction restricts SEATER’s practicality for large datasets or frequent retraining.
@@ 3.2.2 Hybrid clustering
In summary, our greedy and hybrid clustering algorithms address SEATER’s scalability bottleneck by replacing the global capacity-constrained optimization with a parallelizable local heuristic and combining it with precise clustering near the tree leaves. These methods reduce tree construction time by up to two orders of magnitude, down to less than 2% of the original on the largest dataset, while preserving inference efficiency.
@@ 4.1 Datasets
Following the original work [14], we evaluate SEATER on Yelp, Books, and News, and extend the evaluation to a larger-scale dataset, Yambda. Table 2 summarizes the main statistics.
@@ 4.1.1 Yelp
Yelp contains user-business interactions and reviews.222https://business.yelp.com/data/resources/open-dataset/ We use the standard train, validation, and test splits provided in the LightGCN repository [4]. This setup was verified with the SEATER authors to align with their original preprocessing.
@@ 4.1.2 Books
The Books subset of the Amazon review dataset [10].333https://nijianmo.github.io/amazon/ We follow the ComiRec preprocessing pipeline [1], identical to the one used in SEATER’s original implementation.
@@ 4.1.3 News
The small partition of the MIcrosoft News Dataset (MIND) [19].444https://msnews.github.io/ Since the original preprocessing code was unavailable, we reconstructed the setup based on author correspondence. We used the training data of MIND-small and split users by an 8:1:1 ratio into train, validation, and test sets. For each user, positive impressions were concatenated with their interaction history, and sequences were expanded into multiple training instances. During evaluation, we used the first 80% of each sequence as input context and the remaining 20% as targets, consistent with the other datasets.
@@ 4.1.4 Yambda
Yambda [11] is a large-scale open dataset of user-music interactions from Yandex Music.555https://huggingface.co/datasets/yandex/yambda With a substantially larger item pool, it serves as a scalability stress test, clearly exposing tree construction as a bottleneck during training and motivating our proposed algorithms. Following the dataset creators’ preprocessing, we keep only organic interactions and filter out tracks played for less than 50% of their duration. Data is temporally split into 299 days for training, a 30-minute gap, one day for validation, and one day for testing.
@@ 4.2.1 Baselines
We compare SEATER against the baselines reported in the original paper, using the authors’ published results. To verify reproducibility and ensure consistent evaluation, we re-implement three representative baselines: GRU4Rec, SASRec, and BERT4Rec, following the configurations described in the original work [14]. GRU4Rec is implemented via the ComiRec repository.777https://github.com/THUDM/ComiRec/tree/master We implement BERT4Rec using an open-source implementation888https://github.com/jaywonchung/BERT4Rec-VAE-Pytorch, and SASRec based on the SEATER authors’ code. GRU4Rec and SASRec are trained autoregressively by predicting the next item from user history, while BERT4Rec uses a bidirectional Cloze-style objective. Learning rates are selected from {0.001, 0.0001}, with 0.001 performing best in most cases and 0.0001 used for BERT4Rec on the News dataset.
@@ 4.2.2 Tree construction
We build SEATER’s identifier tree using SASRec embeddings trained on the same dataset. These embeddings are only used during tree construction, and then discarded. We use the tree construction implementation provided by the original authors [14], which uses constrained -means [8]. Our greedy variant replaces it with a Faiss-based approximate nearest-neighbor search [3], and the hybrid method switches from greedy to constrained clustering once the number of items per level drops below 2,000, an empirically chosen threshold balancing efficiency and clustering precision.
@@ 4.2.3 Evaluation protocol
Following [14], we evaluate retrieval accuracy using Recall@k, Hit Rate@k (HR@k), and Normalized Discounted Cumulative Gain@k (NDCG@k) for , averaged across test users. To assess training efficiency, we measure total training time (excluding tree construction) and tree-construction time. This allows us to isolate the cost of tree building and highlight its impact as a potential training time bottleneck. It is worth noting that, next to our reported tree-construction and training times, the full SEATER pipeline also includes training SASRec to obtain item embeddings. Following the original work [14], we assume that practitioners already have access to an existing embedding model and therefore exclude SASRec training time from our comparisons.
@@ 5.2 RQ2: The impact of greedy tree construction
Using greedy tree construction in SEATER substantially reduces tree-building time across all datasets, as shown in Table 5 and 1(b). The improvement is particularly pronounced for datasets with larger item catalogs, such as Books and Yambda, where tree-construction time decreases from 54 minutes to 93 seconds and from 82 minutes to 83 seconds, respectively, representing a 35 and 59 speedup. Interestingly, even when excluding tree construction, overall training time also decreases, likely due to faster model convergence. We leave it to future work to investigate the relationship between item identifier structure and convergence dynamics. Despite the use of an approximate (greedy) clustering algorithm, the performance drop across all metrics remains minimal, with the largest observed difference being only in Recall@20 on the Books dataset. To answer RQ2, the proposed greedy clustering method substantially improves training efficiency with a minimal drop in retrieval quality.
@@ 5.3 RQ3: Balancing greedy and complete clustering
| News | NDCG@20 | 0.09160.9e-3 | 0.09591.0e-3 | 0.09501.0e-3 | 0.09572.4e-3 |
| NDCG@50 | 0.12090.8e-3 | 0.12511.2e-3 | 0.12461.1e-3 | 0.12492.4e-3 | |
| Recall@20 | 0.15162.2e-3 | 0.15811.5e-3 | 0.15481.6e-3 | 0.15523.3e-3 | |
| Recall@50 | 0.25242.9e-3 | 0.25813.1e-3 | 0.25761.6e-3 | 0.25552.4e-3 | |
| Hit@20 | 0.41068.6e-3 | 0.42362.6e-3 | 0.41515.0e-3 | 0.41558.2e-3 | |
| Hit@50 | 0.57927.3e-3 | 0.58101.1e-3 | 0.58120.4e-3 | 0.57557.4e-3 | |
| Tree time (s) | – | 190.02 2.55 | 67.08 0.71 | 31.61 0.72 | |
| Training time | 0:11:13 | 1:37:48 | 1:37:18 | 1:35:56 | |
| Yelp | NDCG@20 | 0.04681.8e-3 | 0.05272.0e-3 | 0.05150.6e-3 | 0.05001.0e-3 |
| NDCG@50 | 0.06940.4e-3 | 0.07561.6e-3 | 0.07420.6e-3 | 0.07202.3e-3 | |
| Recall@20 | 0.05971.8e-3 | 0.06542.1e-3 | 0.06390.7e-3 | 0.06161.1e-3 | |
| Recall@50 | 0.11951.6e-3 | 0.12681.6e-3 | 0.12490.7e-3 | 0.12084.8e-3 | |
| Hit@20 | 0.37628.8e-3 | 0.39749.8e-3 | 0.39669.8e-3 | 0.38447.1e-3 | |
| Hit@50 | 0.58593.3e-3 | 0.59401.3e-3 | 0.59005.3e-3 | 0.57918.7e-3 | |
| Tree time (s) | – | 189.35 3.61 | 61.27 0.23 | 31.15 0.34 | |
| Training time | 0:18:42 | 1:10:33 | 1:10:14 | 0:59:42 | |
| Books | NDCG@20 | 0.04080.4e-3 | 0.05930.1e-3 | 0.05740.5e-3 | 0.05660.3e-3 |
| NDCG@50 | 0.05360.5e-3 | 0.07140.2e-3 | 0.06910.5e-3 | 0.06820.4e-3 | |
| Recall@20 | 0.08070.7e-3 | 0.09670.3e-3 | 0.09360.5e-3 | 0.09220.4e-3 | |
| Recall@50 | 0.13111.3e-3 | 0.14420.3e-3 | 0.13980.3e-3 | 0.13740.3e-3 | |
| Hit@20 | 0.16891.2e-3 | 0.20000.8e-3 | 0.19600.9e-3 | 0.19310.3e-3 | |
| Hit@50 | 0.25782.0e-3 | 0.28151.5e-3 | 0.27560.3e-3 | 0.27200.2e-3 | |
| Tree time (s) | – | 3232.80 14.09 | 184.05 0.56 | 93.32 3.92 | |
| Training time | 0:26:49 | 3:54:27 | 3:04:38 | 2:28:32 | |
| Yambda | NDCG@20 | 0.096017.8e-3 | 0.12457.3e-3 | 0.149521.0e-3 | 0.139917.9e-3 |
| NDCG@50 | 0.113821.0e-3 | 0.13708.0e-3 | 0.160120.4e-3 | 0.149317.6e-3 | |
| Recall@20 | 0.100620.0e-3 | 0.11758.0e-3 | 0.146123.6e-3 | 0.135320.1e-3 | |
| Recall@50 | 0.163629.9e-3 | 0.170511.7e-3 | 0.196624.0e-3 | 0.181722.1e-3 | |
| Hit@20 | 0.429834.2e-3 | 0.456618.0e-3 | 0.497528.3e-3 | 0.484127.1e-3 | |
| Hit@50 | 0.540635.3e-3 | 0.548517.6e-3 | 0.579524.5e-3 | 0.563827.2e-3 | |
| Tree time (s) | – | 4907.97 68.16 | 375.72 3.31 | 83.21 2.14 | |
| Training time | 0:47:16 | 5:39:04 | 4:23:58 | 2:37:05 | |
5.4 Ablation Study
To further validate our method, we perform an ablation over the switch depth at which we transition from clustering-based construction to greedy matching. On Books (Table 7), increasing the switch depth improves ranking quality but substantially increases tree construction time, revealing a clear efficiency-effectiveness tradeoff. The fastest construction occurs at depth 1000, while the strongest ranking performance is achieved at depth 50000, at the cost of an order-of-magnitude slower build time. Importantly, gains exhibit diminishing returns: depth 2000 achieves performance comparable to depth 50000 (e.g., nDCG@20 0.0572 vs 0.0585) while reducing construction time substantially (187.07s vs 1452.15s).
This pattern is even more pronounced on Yambda (Table 6), where depth 2000 provides the best ranking metrics overall, while deeper thresholds incur large additional construction costs with little or no additional benefit. Overall, depth 2000 appears to be a robust operating point that balances ranking quality and construction efficiency across datasets.
| Threshold | ||||
| Metric | 100 | 1000 | 2000 | 50000 |
| Time (s) | 79.95 3.28 | 129.99 2.40 | 380.35 4.80 | 2476.85 14.05 |
| nDCG@20 | 0.1507 0.0169 | 0.1495 0.0183 | 0.1616 0.0009 | 0.1509 0.0163 |
| nDCG@50 | 0.1600 0.0168 | 0.1591 0.0173 | 0.1724 0.0009 | 0.1617 0.0161 |
| Hit@20 | 0.4994 0.0246 | 0.4956 0.0268 | 0.5133 0.0015 | 0.4986 0.0250 |
| Hit@50 | 0.5751 0.0223 | 0.5721 0.0215 | 0.5932 0.0038 | 0.5788 0.0234 |
| Recall@20 | 0.1462 0.0178 | 0.1449 0.0206 | 0.1589 0.0009 | 0.1475 0.0182 |
| Recall@50 | 0.1930 0.0200 | 0.1928 0.0201 | 0.2110 0.0014 | 0.1981 0.0199 |
| Threshold | ||||
| Metric | 50 | 1000 | 2000 | 50000 |
| Time (s) | 91.52 0.46 | 90.22 0.61 | 187.07 1.49 | 1452.48 18.28 |
| nDCG@20 | 0.0566 0.0003 | 0.0562 0.0009 | 0.0572 0.0007 | 0.0585 0.0002 |
| nDCG@50 | 0.0682 0.0003 | 0.0678 0.0009 | 0.0689 0.0006 | 0.0703 0.0003 |
| Hit@20 | 0.1931 0.0002 | 0.1922 0.0027 | 0.1958 0.0009 | 0.1974 0.0013 |
| Hit@50 | 0.2720 0.0001 | 0.2713 0.0020 | 0.2754 0.0005 | 0.2774 0.0012 |
| Recall@20 | 0.0922 0.0003 | 0.0917 0.0013 | 0.0938 0.0005 | 0.0949 0.0004 |
| Recall@50 | 0.1374 0.0002 | 0.1368 0.0013 | 0.1394 0.0002 | 0.1413 0.0009 |